Hello, I am in a college Physics course and I am stuck can you please help me, here is the question as it is written on my sheet:
Draw a right triangle so that one angle is 37 degrees and it''s adjacent side is 9 cm, then draw another right triangle so that one angle is 37degress and it's adjacent side is 6.5 cm. Then find the trigonometric ratios of the 37deg angle of each triangle. : sin 37= cos 37= tan 37=......I sure hope this makes sense to you,and I thank you sooo much for your help. I have class in the morning..thanks again

Answer by jamaalprof(6)

it is a special case for 37deg.standard ratio is sin37=3/5,cos37=4/5,tan37=3/4
remember it us useful.in the same way try for 5,12,13